OpenStudy (kendre):

A particle moves along the x-axis with a velocity v given by v(t) = 2t^2 - t + 1. When t = 0, the particle is at x = -3. Find the position along the x axis of the particle at t = 2 A) 7 B) 7 / 3 C) 1 D) 22 E) 3 / 7 I'm 98.8% Sure its B but that 1.2% is getting to me

OpenStudy (trojanpoem):

v = dx/dt dx = 2t^2 - t + 1 dt integrate x from (-3 to x) and t from ( 0 to t ) or (0 to 2) directly x + 3= (2/3)t^3 - (1/2)t^2 + t - 0 x = (2/3)t^3 - (1/2)t^2 + t - 3 at t = 2 x = 7/3

OpenStudy (kendre):

I KNEW IT!!!! Thanks for The assist man

OpenStudy (trojanpoem):

Any time.